Racin' Scooters is introducing a new product and has an expected change in EBIT of $425,000. Racin' Scooters has a 33 percent marginal tax rate. The project will produce $90,000 of depreciation per year. In​ addition, the project will cause the following changes in year​ 1:

What is the​project's free cash flow in year​ 1?

   WITHOUT THE PROJECT   WITH THE PROJECT
Accounts receivable   44,000   62,000
Inventory   65,000   89,000
Accounts payable   75,000   93,000
Answer #1

Free cash flow=net income+ depreciation+ change in net working capital

Net income=ebit-taxes

=425000-(33%*425000)

=284750

depreciation=90000

For change in working capital if there is increase in account receivable or inventory then it is cash outflow and if increase in accounts payable it is cash inflow

change in net working capital=(44000-62000)+(65000-89000)+(93000-75000)

=-24000

Free cash flow=284750+90000-24000=350750